the head of golf club might contains 250 grams of titanium. how many moles is this?
1
Expert's answer
2019-07-01T03:54:10-0400
"n (Ti)=\frac{m}{M} =\frac{250g}{47.87\frac{g}{mol}}=5.2 mol"